During an experiment to determine the molar heat of solution, 2.500 grams of an unknown ionic salt is added to 100.0 grams of water in a coffee cup style calorimeter. The temperature of the water in the calorimeter rises from 20.0 oC to 24.5 oC. Assuming the specfic heat capacity of the solution in the calorimeter is 4.184 J/g oC and the molar mass of the unknown is 120 g/mol, calculate the molar heat of solution of the unknown Use the appropriate sign on your answer and express your answer in kJ/mol. Express your answer to three significant figures.
1
Expert's answer
2016-11-02T15:51:10-0400
The heat released from the solution of the salt can be calculated as Q = m*C*(t1 - t2) = 102.5 g*4.184 J/gC*4.5 C = 1930 J.
The amount of the salt is 2.500 g/120 g/mol = 0.0208 mol.
Thus, the specific heat of the solution is 1930 J/0.0208 mol = 92.8 kJ/mol (the enthalpy of solution is -92.8 kJ/mol).